if tan 0 =(x-1)/4x,then sec 0 -tan 0 = where 0 represents theeta

if tan 0=(x-1)/4x,then sec0-tan0 =


where 0 represents theeta

Grade:12

3 Answers

Chetan Mandayam Nayakar
10 years ago

use the relation sec(theta)=±√(1+tan2(theta))

Vikas TU
10 years ago

From the identity,

sec^2 0 - tan^0 = 1

(sec0 + tan0) (sec0 - tan0) = 1

                            sec0 - tan0  = 1/(sec0 + tan0)

                    

Now from tan 0=(x-1)/4x

                  find the values of cos and sec and put in the above eqn. u will get yr. ansr.!

 

 

 

 

PLZ   Approve!


lakshmi kalyani
10 years ago

Let, secθ+tanθ=p

        secθ-tanθ=1/p then  2ptanθ=1-p2   

there fore substituting the value of tan θ u will get p value and find 1/p

the 1/p is ur result 

PLS APPROVE THANKING U